The measure of an angle is two times the measure of a complementary angle. What is the measure of each angle?
Answer by flame8855(424)

xis the angle and 90-x is the complementary
x= 2 ( 90-x)
x =180 -2x
3x =180
x = 60
90-x=90-60
30 is the complementary